OpenStudy (anonymous):

Solve 3x2 + 2x + 7 = 0. Round solutions to the nearest hundredth

OpenStudy (anonymous):

3x² + 2x + 7 = 0 using the abc-formula: a=3, b=2, c=7; D=b²-4ac=4-4*3*7=4-81=-77 So x is no Real Number.